OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1).

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

the given line is :y=-3x+7 so its slope is -3 the slope of the perpendicular is 1/3 so its equation is y=(1/3)*x+n now substitute (6;-1) in it: -1=2+n=>n=-3 so:y=(1/3)*x-3